### v1

# LLM Knows Geometry Better than Algebra: Numerical Understanding of LLM-Based Agents in A Trading Arena

Abstract
Recent advancements in large language models (LLMs) have significantly improved performance in natural language processing tasks. However, their ability to generalize to dynamic, unseen tasks, particularly in numerical reasoning, remains a challenge. Existing benchmarks mainly evaluate LLMs on problems with predefined optimal solutions, which may not align with real-world scenarios where clear answers are absent. To bridge this gap, we design the Agent Trading Arena, a virtual numerical game simulating complex economic systems through zero-sum games, where agents invest in stock portfolios. Our experiments reveal that LLMs, including GPT-4o, struggle with algebraic reasoning when dealing with plain-text stock data, often focusing on local details rather than global trends. In contrast, LLMs perform significantly better with geometric reasoning when presented with visual data, such as scatter plots or K-line charts, suggesting that visual representations enhance numerical reasoning. This capability is further improved by incorporating the reflection module, which aids in the analysis and interpretation of complex data. We validate our findings on NASDAQ Stock dataset, where LLMs demonstrate stronger reasoning with visual data compared to text. Our code and data are publicly available at https://github.com/wekjsdvnm/Agent-Trading-Arena.git.
LLM Knows Geometry Better than Algebra:
Numerical Understanding of LLM-Based Agents in A Trading Arena
Tianmi Ma1, Jiawei Du3,4,, Wenxin Huang2, ✉, Wenjie Wang5, Liang Xie6, Xian Zhong1, ✉, and Joey Tianyi Zhou3,4 1 Hubei Key Laboratory of Transportation Internet of Things, Wuhan University of Technology 2 Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University 3 Centre for Frontier AI Research, Agency for Science, Technology and Research, Singapore 4 Institute of High Performance Computing, Agency for Science, Technology and Research, Singapore 5 School of Computing, National University of Singapore  6 School of Science, Wuhan University of Technology
footnotetext: Email:wenxinhuang_wh@163.com, zhongx@whut.edu.cn. denotes the equal contribution. ✉ represents the corresponding author.
1 Introduction
Recent advancements in large language models (LLMs) have demonstrated exceptional proficiency across various domains, achieving state-of-the-art performance in natural language processing (NLP) tasks such as translation (Koshkin et al., 2024), summarization (Yang et al., 2023), and reasoning (Kalyanpur et al., 2024; Qiao et al., 2023). While LLMs excel in language-based tasks, further progress in numerical and geometric reasoning is essential to tackling complex, interdisciplinary challenges, particularly in fields like finance and scientific research. Benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) have been developed to assess and improve the mathematical problem-solving abilities of LLMs through structured datasets and standardized evaluation protocols. These benchmarks not only advance LLM development but also represent a crucial step in enhancing the ability of artificial intelligence systems to solve mathematical reasoning problems.
Figure 1: Illustration and Performance Comparison of Textual and Visual (K-Line Chart) Numerical Inputs. Top: The left figure presents numerical data in K-line chart format, while the right figure shows textual numerical inputs. Bottom: LLMs with visual inputs (orange) significantly outperform those with textual inputs (blue) in 40-day yield trends, demonstrating the advantage of visual data in financial decision-making.
In response, recent updates to LLMs have focused on improving their mathematical abilities, with models (Ahn et al., 2024; Romera-Paredes et al., 2024) achieving new records on math-focused benchmarks. However, many of these benchmarks (Taylor et al., 2022; Naveed et al., 2023) primarily evaluate performance on problems familiar to the models, closely resembling those encountered during training and often following standard formats, solution strategies, or recurring mathematical patterns. This raises concerns about whether LLMs genuinely possess reasoning abilities for novel numerical data or whether their success is largely due to memorization and pattern recognition. These limitations underscore the need for alternative evaluation paradigms that assess LLMs’ ability to generalize mathematical principles to unseen scenarios.
To address these limitations, we introduce the Agent Trading Arena, a virtual numerical game generating diverse numerical data through agent interactions in a zero-sum stock market environment. In this system, LLM-based agents make trading decisions based on historical stock prices, dynamically determined by agents’ bidding activities. Unlike static benchmarks (Naveed et al., 2023) with predefined strategies, our simulation forces agents to adapt to evolving market conditions. Any temporarily optimal strategy is quickly countered, fostering an environment that rewards adaptability. By mimicking competitive trading scenarios where outcomes depend on real-time agent decisions, we create a controlled yet dynamic environment that tests LLMs’ ability to infer hidden patterns and generalize numerical laws without relying on memorization or prior knowledge.
Our experiments reveal that LLMs struggle with textual numerical data, often focusing on absolute values rather than capturing percentage changes and relations between data points, akin to algebraic reasoning. Additionally, LLMs tend to overemphasize recent data while overlooking earlier information, even when explicitly highlighted in the prompt. These limitations suggest that LLMs lack the ability to abstract numerical information into higher-level representations, a crucial skill for generalizing beyond explicit values.
In contrast, LLMs perform better when presented with numerical data in visualized formats, such as scatter plots, line charts, and bar graphs, which involve geometric reasoning. Our experiments show that LLMs processing visual numerical data consistently outperform those processing textual numerical data. As shown in Figure 1, visual numerical data enables LLMs to achieve significantly higher return rates in the Agent Trading Arena, underscoring the advantages of structured visual representations.
Incorporating the reflection module (Zhou et al., 2022; Shinn et al., 2023) further emphasizes the performance gap between textual and visual data, with the module proving particularly effective for visual representations. By leveraging structured visualizations, the reflection module enhances the model’s ability to reason through complex data relations, leading to more accurate and strategic decision-making.
The Agent Trading Arena provides a valuable framework for advancing LLM generalization and adaptability, with broad applications in fields like finance, healthcare, and scientific research. To further validate our findings, we extend our evaluation to real-world financial data, such as NASDAQ Stock dataset, which includes historical trading data. The consistent results suggest an important contribution: LLMs exhibit stronger reasoning with visual geometric data compared to textual numerical data, demonstrating the value of visual representations in enhancing LLM performance.
2 Related Works
2.1 Mathematical Benchmarks for LLMs
Math word problems (MWPs) have been widely studied, leading to the development of various benchmarks for evaluating models’ mathematical reasoning and problem-solving abilities. Early datasets, such as MAWPS (Koncel-Kedziorski et al., 2016), standardized existing problems to facilitate consistent evaluation. Math23K (Wang et al., 2017) introduced a large-scale collection of Chinese arithmetic problems that require structured equation solving. To increase diversity, benchmarks like ASDiv (Miao et al., 2020) and SVAMP (Patel et al., 2021) provide richer annotations and a broader range of problem types. More recent benchmarks, including GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), focus on multi-step reasoning and advanced mathematical concepts, broadening the scope of evaluation. Additionally, MathQA-Python (Austin et al., 2021), a Python variant of MathQA (Amini et al., 2019), emphasizes programmatic reasoning, while MGSM (Shi et al., 2023) extends these benchmarks to multilingual contexts. Despite these advancements, current models (Lu et al., 2023b, a) primarily rely on memory-based answering strategies learned through extensive training, rather than demonstrating true mathematical reasoning.
2.2 LLMs for Enhanced Mathematical Reasoning
Building on these benchmarks, LLMs have advanced mathematical problem-solving by incorporating specialized datasets into their training. Models such as Galactica (Taylor et al., 2022), PaLM-2 (Anil et al., 2023), Minerva (Lewkowycz et al., 2022), and LLaMA-2 (Touvron et al., 2023) leverage extensive datasets during pre-training, improving their mathematical reasoning abilities and understanding of complex concepts. Fine-tuned models like MetaMath (Yu et al., 2024), MAmmoTH (Yue et al., 2024), and WizardMath (Luo et al., 2023) are specifically tailored for mathematical tasks. These models undergo domain-specific fine-tuning with carefully curated datasets, enabling them to tackle advanced reasoning challenges and significantly enhance performance on specialized benchmarks.
However, existing methods (Naveed et al., 2023; Yang et al., 2024a) often rely on large training datasets, raising concerns about the true reasoning capabilities of LLMs. While these methods primarily assess performance on established problem types, the heavy reliance on extensive training data suggests that models may achieve high performance through memorization and pattern recognition, rather than genuine reasoning. Consequently, alternative evaluation paradigms are needed to more accurately assess LLMs’ ability to generalize mathematical principles to novel scenarios.
Figure 2: Stock Trading Workflow in Agent Trading Arena. Top: Workflow of a trading day, including preparation, trading, and post-trading reflection. Agents discuss insights in the chat pool, analyze market trends, execute trades, and refine strategies based on performance. Bottom: Example of agents’ interactions in the chat pool and dynamic strategy updates.
3 Proposed Method
To mitigate the influence of human prior knowledge and memory, we designed a closed-loop economic system (Guo et al., 2024) called the Agent Trading Arena, a zero-sum game simulating complex, quantitative real-world scenarios. The simulation workflow is illustrated in Figure 2 and further detailed in Appendix A. In the Agent Trading Arena, agents can invest in assets, earn dividends from holding assets, and pay daily expenses using virtual currency. The agent with the highest total return wins the game.
3.1 Agent Trading Arena
Structure of Agent Trading Arena.
To eliminate external knowledge biases, asset prices are determined by a bid-ask system, reflecting the prices at which buyers and sellers are willing to transact. The system evolves solely based on agents’ actions and interactions, without external influences. This design ensures that the outcomes of agents’ actions are not immediately apparent but unfold gradually, influenced by other agents’ decisions.
To encourage active participation, a dividend mechanism is introduced. There are two primary sources of income in this system: capital gains from asset price differentials and dividends from holding assets. Dividends for each asset are distributed according to a predefined ratio, serving as an implicit anchor for asset prices. Agents holding more low-cost assets receive higher dividends. To prevent passive asset holding until the end of the game, agents must pay a daily capital cost proportional to their total wealth. These expenses are offset by asset dividends, and only agents with sufficient low-cost assets can cover costs. Under the pressure of significant daily expenses, agents must act swiftly and strategically, triggering frequent trades and price fluctuations to stimulate market activity. This dynamic mechanism ensures fairness in the zero-sum game while preventing agents from relying on fixed strategies to find optimal solutions.
Agents Learn and Compete in Arena.
The zero-sum game structure is crucial to eliminating the possibility of a universally optimal strategy. In fixed scenarios with a static optimal solution, agents could rely on predefined rules or memory-based approaches, bypassing adaptive decision-making. The zero-sum game ensures that there is no universally correct solution, with outcomes evolving dynamically based on agent interactions and competition. This design forces agents to continually adapt, learn from feedback, and develop context-dependent strategies, promoting deeper environmental exploration and preventing reliance on static or memory-driven solutions.
In the Agent Trading Arena, agents are unaware of implicit rules, except for the objective to maximize their virtual wealth throughout the simulation. To win this zero-sum game, agents must effectively learn from experience, decipher hidden game rules, and develop strategies to counter competitors. This requires the ability to comprehend numerical feedback, formulate enduring strategies, and make informed decisions. Unlike other mathematical reasoning problems, the results of their actions unfold gradually and dynamically. Moreover, agents are easily misled by erroneous information from competitors, hindering their ability to discern strategic cues from competitors’ textual data. Importantly, agents remain unaware of these implicit rules, so applying real-world knowledge does not benefit their performance. Therefore, agents must rely on experiential learning to decipher the hidden game rules and ultimately achieve victory.
3.2 Types of Numerical Data Input
Limitations of Textual Numerical Data.
In the Agent Trading Arena, the generated stock data is stored in numerical format. When used directly as input to an LLM, the models often struggle to interpret numerical data accurately or make sound decisions. To mitigate this, we convert the data into textual formats (Hu et al., 2024; Zhang et al., 2024), enhancing semantic features and clarifying output requirements to improve the models’ understanding. During interactions, the LLMs process stock prices, trading volumes, and market indices presented as textual numerical data.
Figure 3: Textual and Visual Representations of Corresponding Inputs and Outputs. The left images display the agent’s Buy and Sell trading records, daily trade prices, and K-line charts for three stocks. The output from visual inputs (bottom right) captures overall stock trends and long-term behavior, while the output from textual inputs (top right) focuses on specific current prices.
However, this textual approach reveals significant limitations. While the data is presented clearly, LLMs tend to focus excessively on specific values rather than identifying long-term trends or global patterns. They also struggle with understanding correlative relations and percentage changes, limiting their ability to assess differences and identify connections between data points. When analyzing time-series data with complex patterns, LLMs often fixate on individual data points, overlooking overarching relations. This issue is evident in the analysis output in the top-right corner of Figure 3, where LLMs’ focus on individual values impedes their ability to generalize, reducing their capacity to extract meaningful global insights.
Additionally, LLMs often overemphasize recent data while undervaluing historical information, even when prompted to consider its importance. This prevents them from effectively integrating past data and recognizing long-term patterns, complicating their understanding of numerical relations and trends. These challenges highlight the need for improved mechanisms to process numerical relations, identify global trends, and derive deeper insights from textual numerical data.
Potential of Visual Numerical Data.
Since textual numerical data often leads LLMs to focus on local details while neglecting broader relations, we investigated whether visual representations, such as scatter plots, line charts, and bar charts, could help LLMs better understand overall trends, similar to human reasoning. Thus, we transition from textual numerical data inputs to visualized formats  (Yang et al., 2025). As demonstrated in the bottom-right corner of Figure 3, visual representations enable LLMs to more effectively grasp global trends, patterns, and relations that are often difficult to discern from textual numerical data alone.
These findings highlight the advantages of structured, visual numerical data, indicating that this format allows LLMs to more intuitively and comprehensively understand complex data, better capturing overall fluctuations, whereas text tends to focus on local details. By combining visualization and textual representations, LLMs not only overcome the challenges of relations in time-series data but also demonstrate better performance in identifying long-term trends and global patterns, while still attending to local details.
3.3 Reflection Module
We propose a strategy distillation method, illustrated in Figure 4, that delivers real-time feedback to LLMs by analyzing both descriptive textual and visual numerical data. This enables the generation of new strategies and optimization of action plans. The approach allows agents to evaluate their results, refine strategies, and adapt continuously based on feedback. The process begins with assessing the day’s trajectory memory and associated strategies using an evaluation function. The strategic generation process leverages contrastive analysis of peak and nadir performers from the evaluation phase, creating bidirectional learning signals that inform subsequent iterations. This iterative cycle ensures continuous strategy evolution, fostering sustained improvement in decision-making.
Figure 4: Design of the Reflection Module. The process evaluates daily trajectory memory and strategies (top right), then generates new strategies (center) based on evaluation, environmental feedback (bottom right), and feedback from the 5 top- and bottom-performing strategies. Stock visualization (bottom left) enhances reflection, driving continuous improvement.
The reflection module plays a crucial role in refining strategies by offering real-time feedback. It analyzes both descriptive textual and visual numerical data to generate new strategies and optimize action plans. Within the Agent Trading Arena, the reflection module is triggered regularly to consolidate daily trading records and evaluate the effectiveness of strategies, refining both successful and unsuccessful experiences to guide future decisions. Ineffective strategies are stored in a strategy library for future reference, allowing agents to review and learn from past experiences. Further details can be found in Appendix A.
ID Name Year Cash Occupation
1 Amy 1 100,000 AI Researcher
2 Bruce 2 100,000 Lawyer
3 Charles 1 100,000 Doctor
4 David 3 100,000 Engineer
5 Ella 2 100,000 Teacher
6 Frank 5 100,000 Entrepreneur
7 Grace 4 100,000 Accountant
8 Hank 2 100,000 Architect
9 Ivy 3 100,000 Marketing Manager
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
Table 1: Agent Details. “Year” is the investment duration, and “Cash” is the initial capital.
ID Name DPS Historical Close Prices Quantity
1 A 22 454.17, $\dots$, 445.60 1,200
2 B 23 354.17, $\dots$, 465.80 1,000
3 C 25 500.47, $\dots$, 440.60 1,600
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
Table 2: Stock Details. “DPS” is the Dividend Per Share, and “Quantity” is the initial share count.
4 Experimental Results
4.1 Experimental Setup
Datasets.
To evaluate the ability of LLMs to analyze and process data, we developed the Agent Trading Arena, a controlled environment that isolates external factors. The system’s workflow is illustrated in Figure 2. Within this environment, agents discuss stock market trends, analyze stock data, and engage in trading activities. Each agent can execute multiple trades per day and reflect on all trades at the end of each trading session.The Agent Trading Arena allows for adjustable numbers of agents and stocks; in our experiment, we deployed at least nine agents and three stocks. All agents were provided with the same initial capital to ensure identical starting conditions. Detailed information about each agent is presented in Table 1, and stock-specific data is available in Table 2. To further validate our findings, we selected a subset of NASDAQ Stock dataset for portfolio investment. For more details, refer to Appendix B.
Evaluation Metrics.
Each agent was assigned varying capital based on their roles, and performance was evaluated using the following metrics:
1) Total Return (TR): Measures the overall performance of the strategy, calculated as: $\mathrm{TR}={(C_{1}-C_{0})}/{C_{0}}$, where $C_{0}$ is the initial asset value and $C_{1}$ is the final asset value.
2) Win Rate (WR): Represents the proportion of winning trades, calculated as: $\mathrm{WR}={N_{w}}/{N_{t}}$, where $N_{w}$ is the number of winning trades and $N_{t}$ is the total number of trades.
3) Sharpe Ratio (SR): Evaluates investment returns relative to risk, calculated as: $\mathrm{SR}={(R_{p}-R_{f})}/{\sigma_{p}}$, where $R_{p}$ is the mean daily return, $\sigma_{p}$ is the standard deviation of daily returns, and $R_{f}$ is the risk-free return, set to 0 as defined in SocioDojo (Cheng and Chin, 2024).
4) Mean Daily Return (Mean): Indicates the average return per day during the trading period.
5) Standard Deviation of Daily Returns (Std): Reflects the volatility of daily returns, indicating the risk associated with the strategy.
4.2 Comparative Experiments
We conducted experiments to assess the real-time data analysis and reasoning capabilities of LLM-based agents, focusing on how textual and visual representations influence decision-making. First, we explored the impact of textual and visual representations in dynamic environments. Next, we incorporated the reflection module to enhance agents’ reasoning and data interpretation, examining how reflective reasoning influences decision-making. Finally, we validated the model’s effectiveness through stock investment simulations on NASDAQ Stock dataset, assessing agents’ adaptability and decision-making in real-world scenarios.
LLMs Textual Visual TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
w/o Reflection LLaMa-3 (Dubey et al., 2024) 4.0934 0.6857 2.2700 62.5001 0.1963
DeepSeek (Liu et al., 2024) 27.3078 2.2478 2.7553 92.2078 1.5303
Qwen-2.5 (Yang et al., 2024b) 30.3740 2.6795 1.4588 93.7500 1.7025
Gemini-1.5 (Reid et al., 2024) 14.3193 1.2210 2.7776 83.8384 1.7293
19.0389 1.6038 1.1654 90.9091 1.3761
23.9649 1.9809 1.3342 100.0000 1.4847
GPT-4o (Hurst et al., 2024) 13.0369 1.1386 1.9285 54.5455 0.5904
17.0661 1.4539 1.5040 72.7272 0.9668
26.1806 2.1574 2.0578 90.9091 1.0484
w/ Reflection LLaMa-3 (Dubey et al., 2024) 10.2458 1.6669 2.4034 66.6667 0.6936
DeepSeek (Liu et al., 2024) 30.6238 2.4696 1.5121 100.0000 1.6332
Qwen-2.5 (Yang et al., 2024b) 38.9113 3.3244 1.0726 100.0000 3.0994
Gemini-1.5 (Reid et al., 2024) 29.4511 2.5185 3.0116 93.5064 2.2100
37.0054 2.9111 1.2349 100.0000 2.3574
41.3264 3.1946 0.1588 100.0000 20.1128
GPT-4o (Hurst et al., 2024) 33.6508 2.6713 2.1711 98.7013 2.1417
35.7622 2.8206 0.6782 100.0000 4.1590
47.6851 3.6095 0.5327 100.0000 6.7765
Table 3: Performance Comparison Using Textual and Visual Data w/o and w/ Reflection. Evaluation of agent portfolios show that visual approaches outperformed textual ones, with the best results achieved by combining both. LLMs demonstrated a significant advantage in TR and SR, with visual inputs yielding greater improvements than textual inputs. The best and second-best results are highlighted in bold and underlined.
Trials with Textual or Visual Input.
To enhance LLMs’ understanding of complex data in the Agent Trading Arena, we transitioned from textual numerical inputs to visualized formats, including scatter plots, line charts, and bar charts. Three types of visualizations were used: daily K-line charts, transaction histories, and quantities traded by each agent. The experimental setup for the visualization group is shown in Figure 3 and detailed in Appendix C. In the Arena, LLMs without image input capabilities received only text input and did not perform reflection. For image-enabled LLMs, the first agent received only visual input, the second received both textual and visual input, and the others received only textual input. None of the agents had reflection capabilities. For details on the selected LLMs, please refer to Appendix E.
We conducted experiments across different LLMs and the results for Gemini-1.5 (Reid et al., 2024) and GPT-4o (Hurst et al., 2024) in Table 3 show that agents with visual numerical input outperformed those with textual input alone. Agents receiving both textual and visual input achieved the best performance. This combination enables agents to focus on local details while also understanding overall trends, leading to optimal performance. This suggests that LLMs exhibit better reasoning skills with visual geometric data rather than textual numerical data, highlighting a clear distinction in their strengths in geometry and algebra.
Strategy TR $\uparrow$ SR $\uparrow$
MACD (Chong and Ng, 2008) 7.18 0.173
StockFormer (Gao et al., 2023) 9.05 0.073
TimesNet (Wu et al., 2023) 11.74 0.203
GPT-4o + Textual 8.69 0.167
GPT-4o + Visual 9.91 0.195
GPT-4o + Textual + Visual (Ours) 12.23 0.291
Table 4: Reproduced Backtesting Performance on NASDAQ Stock Dataset. The agent with textual and visual input outperformed the NASDAQ-100 by 53.97% in SR during the same period.
LLMs T V TR $\uparrow$ SR $\uparrow$
LLaMa-3 100.00 100.00
DeepSeek + 35.03 + 45.63
Gemini-1.5 + 10.99 + 14.93
+ 11.49 + 29.47
GPT-4o + 10.39 + 28.29
+ 17.18 + 40.77
Table 5: Performance Comparison in Trading Decisions w/o Reflection, Competing Pairwise with LLaMa-3. DeepSeek may possess unique strengths or optimizations that allow it to better adapt to the task’s complexities when competing with other models.
LLMs TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
LLaMa-3 (Dubey et al., 2024) 100.0000 100.0000 100.0000 100.0000 100.0000
Gemini-1.5 (Reid et al., 2024) + 3.8631 + 5.9249 + 12.0426 - 4.0010 + 2.2920
DeepSeek (Liu et al., 2024) + 3.1317 + 3.1309 + 1.7169 + 9.8039 + 1.7945
Qwen-2.5 (Yang et al., 2024b) + 7.4578 + 7.3907 - 2.4801 + 50.0000 + 8.0524
GPT-4o (Hurst et al., 2024) + 17.7636 + 16.9041 + 2.4136 + 37.9310 + 14.2775
Table 6: Performance Comparison in Trading Decisions w/ Reflection, Competing Pairwise with LLaMa-3 Using Textual Data. The results show that in competition with reflection using textual numerical data, GPT-4o and Qwen-2.5 outperformed other models.
Trials with Reflection and Textual or Visual Input.
Building upon the previous experiments, we incorporated the reflection module to investigate its impact on the experimental outcomes. By introducing this module, we aimed to determine whether reflective processing could improve strategies and adaptability in complex financial scenarios.
Our experiments with diverse LLMs demonstrate that reflective agents consistently outperformed non-reflective agents in stock trading. As shown in Table 3, text-only LLMs with reflection modules achieved enhanced returns, with GPT-4o (Hurst et al., 2024) reaching 33.65% total return using textual input alone, which increased to 47.70% with visual data. Combining both input types resulted in a 41.74% higher return compared to using textual input alone. Compared to agents without the reflection module, including the reflection module significantly enhanced each agent’s stock investment performance. LLMs with the reflection module showed the greatest improvement in agents with visual inputs, compared to those with textual inputs, in both the Total Return and Sharpe Ratio. This indicates that incorporating reflection further widens the gap between LLMs’ geometric reasoning with visual data and algebraic reasoning with textual numerical data, highlighting their strengths in both areas.
Simulation with NASDAQ Stock dataset.
To further validate our findings, we conducted a two-month investment simulation on NASDAQ Stock dataset, starting with an initial capital of 100,000 units for portfolio investment. NASDAQ Stock dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays. Detailed information about the models used is provided in Appendix B.
The results in Table 4 show that, despite StockFormer (Gao et al., 2023) and TimesNet (Wu et al., 2023) benefiting from longer training periods and larger datasets, they underperformed compared to our model. Notably, our model achieved superior results without additional training, relying solely on historical stock data for decision-making. The Sharpe ratios for Apple, NASDAQ-100, and S&P 500 during the same period were 0.097, 0.189, and 0.205, respectively. The agent with textual and visual input outperformed the NASDAQ-100 and S&P 500 in Sharpe Ratio by 53.97% and 41.95%, respectively. Moreover, only visual input outperformed textual input, further highlighting LLMs’ stronger geometric reasoning abilities.
4.3 Ablation Experiments
Impact of Modality on LLM Competitiveness.
We employed a relative evaluation method for this experiment. The first and second agents used various LLMs in textual and visual settings, respectively, while the remaining agents were based on LLaMa-3 (Dubey et al., 2024) as the baseline. This setup aimed to explore the impact of different agents and modalities on LLM performance. The results are shown in Table 5. The findings indicate that DeepSeek (Liu et al., 2024) exhibited stronger competitive performance across different LLM environments, suggesting unique strengths or optimizations that enable it to adapt more effectively to the task’s complexities.
Competition Among Different LLMs with Reflection.
To investigate the role of reflection in competition dynamics using textual numerical data between LLMs, we conducted an ablation study in the Agent Trading Arena. Using a relative evaluation approach, the first agent employed various LLMs, while eight LLaMa-3-based agents served as the baseline for a fair comparison. This setup effectively isolated the influence of each LLM on performance. As shown in Table 6, in competition with reflection, GPT-4o (Hurst et al., 2024) and Qwen-2.5 (Yang et al., 2024b) outperformed other models, consistent with the findings in Table 3.
5 Conclusion
In this paper, we introduced the Agent Trading Arena, a zero-sum game designed to simulate complex economic systems and evaluate LLMs on numerical reasoning tasks. Our results show that while LLMs struggle with plain-text numerical data (algebraic reasoning), their performance significantly improves when presented with visual data (geometric reasoning). This highlights the advantage of visual representations in supporting numerical reasoning, particularly in complex scenarios. The integration of a reflection module further enhances model performance, allowing LLMs to analyze better and interpret data. We validated these findings on NASDAQ Stock dataset, demonstrating that LLMs excel in visual geometric reasoning tasks, suggesting that LLMs may perform better with visual numerical data than with textual numerical data. Overall, our work offers insights into the strengths and limitations of LLMs in dynamic numerical reasoning tasks, particularly in the context of geometry vs. algebra, and sets the foundation for future research on improving their performance in real-world, interdisciplinary challenges.
Limitations
This study evaluates LLMs within a virtual stock trading environment, focusing on their performance in visual geometric reasoning. While this controlled setting limits the generalizability to other domains, it provides valuable insights into LLMs’ capabilities. The reliance on high-quality visualizations, reflection modules, and substantial computational resources may restrict applicability in resource-constrained environments. Future research can address these limitations by broadening the scope to include diverse reasoning tasks, optimizing computational requirements, and exploring alternative modalities for more robust and generalizable assessments.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62301213 and 62271361, the Hubei Provincial Key Research and Development Program under Grant 2024BAB039, and the Open Project Funding of the Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University under Grant 2024BDIAA01.
Appendix A Agent Trading Arena
A.1 Agent Trading Arena Details
As illustrated in Figure 5, each agent’s workflow integrates LLMs for chat pool interactions, stock analysis, decision-making, and reflection. In the stock analysis and decision-making modules, all outputs are validated for consistency with both common sense and operational requirements before execution.
Figure 5: Agent Workflow Components: Share Trading involves stock analysis, decision validation, and trade execution; Environment and Memory manage memory and process trade orders; and Reflection focuses on strategy assessment and refinement based on feedback.
A.1.1 Action Decision-Making
Action generation follows the LLM framework. The agent responsible for generating actions receives corresponding prompts via SQLite. Based on these prompts and specified output formats, the agent decides whether to buy, sell, or hold stocks. The action generation process is outlined in Equation 1, with input prompts shown in Figure 6 and the corresponding outputs displayed in the adjacent figure.
$\begin{cases}A_{\mathrm{date}+1}^{t}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{t},S_{\mathrm{date}}\right),&\text{if~{}}t\text{~{}is Iters},\\ A_{\mathrm{date}}^{t+1}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{t},S_{\mathrm{date}}\right),&\text{otherwise},\end{cases}$ (1)
where $\mathrm{Ins}$ represents the environment introduction, $Z_{\mathrm{date}}^{t}$ denotes the memory of the stock transaction on day $\mathrm{date}$ retrieved from the database, and $S_{\mathrm{date}}$ is the strategy for day $\mathrm{date}$ generated via reflection.
Figure 6: Inputs and Outputs in Action Decision-Making.
A.1.2 Environmental Interaction
To isolate external influences, we created a virtual sandbox environment where each agent is assigned a unique ID, and their actions affect the environment. The function $\phi$ facilitates environmental interactions, as shown in Algorithm 1, where “OPS” retrieves agent actions, “date” refers to the trading date, and “Z” represents the memory used for interaction with the environment. Through $\phi$, each agent’s actions, such as buying or selling stocks, determine the current stock price and update the trading platform, including stock prices and available shares. Stock prices are independent of external factors and are influenced solely by the sandbox’s internal dynamics. The stock price is updated with each transaction according to the following formula:
$\mathrm{Price}_{\mathrm{curr}}$ $=\delta\left(Q,F,\mathrm{Price}_{\mathrm{curr}},\mathrm{Price}_{\mathrm{deal}}\right)$ (2)
$=\frac{\mathrm{Price}_{\mathrm{deal}}\cdot Q\cdot F+\mathrm{Price}_{\mathrm{curr}}\cdot Q_{\mathrm{total}}}{Q\cdot F+Q_{\mathrm{total}}},$
where $Q$ is the quantity of stock traded, $F$ is the fluctuation constant, $\mathrm{Price}_{\mathrm{curr}}$ is the current stock price, $Q_{\mathrm{total}}$ is the total number of shares available, and $\mathrm{Price}_{\mathrm{deal}}$ is the price at which the trade occurs.
Input: OPS: Function to retrieve agents’ actions, date: The current trading date, Z: Memory used for interaction with the environment
Output: Z
for $P\in\text{Persons}$ do
       $A\leftarrow\text{OPS}(t,P)$;
       $O,N,Q,\text{Price}_{\text{deal}}\leftarrow\text{Extract}(A)$;
       $\text{Price}_{\text{curr}},Q_{\text{total}}\leftarrow\text{Stocks}(N)$;
       $\text{Price}_{\text{curr}}\leftarrow\delta(Q,F,\text{Price}_{\text{curr}},\text{Price}_{\text{deal}})$;
      if $O=$ “buy” then
             $\text{Cash}\leftarrow\text{Price}_{\text{curr}}\cdot Q$;
             if $<P.\text{Cash}$ Cash then
                   $Z\leftarrow\text{SubmitOrder}(O,N,t,\text{Price}_{\text{curr}},Q)$;
      if $O=$ “sell” then
             $\text{Hold}\leftarrow P(N)$;
[…]
                   $Q_{N}\leftarrow\text{Hold}[\text{``Q''}]-Q$;
                   if $Q_{N}>0$ then
                         $Z\leftarrow\text{SubmitOrder}(O,N,t,\text{Price}_{\text{curr}},Q)$;
[…]
return Z;
Algorithm 1 Environmental Interaction
The function $\phi$ executes trading orders and updates the stock price in real-time based on the agent’s actions. To prevent excessive volatility and mitigate risk during trading cycles, a daily price fluctuation cap is enforced. Before executing any transaction, each agent evaluates its available funds and refrains from proceeding if insufficient capital is available.
A.1.3 Memory
The superior performance of LLM-based agents arises from the extensive internal knowledge acquired during pre-training. The large number of parameters in LLMs enables the retrieval of diverse information and supports logical and inferential reasoning. To further enhance knowledge retrieval across various tasks, we incorporate a memory module that empowers LLM-based agents with self-improvement capabilities. This memory module facilitates strategy reflection through time-series feedback. Unlike qualitative tasks, quantitative feedback evolves incrementally with subtle differences, presenting a challenge for the generalization of existing LLM-based agents.
To minimize the influence of pre-existing knowledge, we assign specific roles to each agent. The memory module records sensory inputs and accumulates valuable experiences based on immediate feedback following actions. These experiences are stored in a database for future reference. The interaction history between an agent’s behavior and the environment constitutes short-term memory, enabling the agent to retain recent events. The historical trajectory is defined as:
$M_{\mathrm{date}}^{t}=\zeta\left(I_{\mathrm{date}}^{t},\mathrm{Out}_{\mathrm{date}}^{t}\right),$ (3)
where $\zeta$ processes key information, $I_{\mathrm{date}}^{t}$ represents the input prompt, and $\mathrm{Out}_{\mathrm{date}}^{t}$ denotes the resulting output.
The day’s trajectory constitutes short-term memory, which is expressed as:
$Z_{\mathrm{date}}^{t}=\left(M_{\mathrm{date}}^{0},M_{\mathrm{date}}^{1},\dots,M_{\mathrm{date}}^{t}\right).$ (4)
The process for updating short-term memory is given by:
$Z_{\mathrm{date}}^{t+1}=M_{\mathrm{date}}^{t}\cup Z_{\mathrm{date}}^{t},\quad t\in\{0,1,\dots,T\},$ (5)
where $Z_{\mathrm{date}}^{t}$ represents the short-term memory, and $T$ denotes the maximum number of iterations.
The reflection model serves as long-term memory, enabling self-reflection and the consolidation of knowledge.
Figure 7: Visualization of Stock Inputs and Corresponding Trading Strategy Outputs.
A.1.4 Reflection
We propose a strategy distillation method that transforms quantitative results into descriptive text, which is then used as prompts for LLMs. This approach aids in the analysis of results and the generation of actionable, qualitative summaries, enabling LLMs to derive new strategies. These strategies are implemented, monitored, and evaluated over time, while underperforming strategies are archived for future review.
Initially, we evaluate the day’s trajectory memory and associated strategies. The evaluation function is defined as:
$E_{\mathrm{date}}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{T},S_{\mathrm{date}}\right),$ (6)
where $\Psi$ represents the evaluation function based on LLMs, $\mathrm{Ins}$ contains the agent role descriptions and output requirements, and $Z_{\mathrm{date}}^{T}$ denotes the memory for the given day.
Based on the evaluation results and previous strategies, we generate the latest strategy. Past strategies are stored in a library and scored. For the new strategy, we select the top five best-performing and the bottom five worst-performing strategies to provide both positive and negative feedback. The strategy update formula is:
$S_{\mathrm{date}+1}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{T},E_{\mathrm{date}},S_{\mathrm{date}}\right),$ (7)
where $E_{\mathrm{date}}$ represents the evaluation results for the day, and $S_{\mathrm{date}}$ is the strategy for that day.
Long-term memory is generated through reflection, as represented by:
$Z^{\prime}_{\mathrm{date}}=\left(S_{0},S_{1},\dots,S_{\mathrm{date}}\right).$ (8)
The process for updating long-term memory is defined as follows:
$Z^{\prime}_{\mathrm{date}+1}=S_{\mathrm{date}}\cup Z^{\prime}_{\mathrm{date}},$ (9)
$\mathrm{date}\in\{0,1,\dots,\mathrm{DAYS}\}.$
where $Z^{\prime}_{\mathrm{date}}$ represents the long-term memory, and $\mathrm{DAYS}$ denotes the maximum number of days.
Together, short-term and long-term memory provide essential context for the agents. Success in this environment depends on their ability to understand the game rules and develop strategies that outmaneuver competitors. Agents continuously refine their strategies based on incremental quantitative feedback, adjusting their actions to align with long-term objectives.
Figure 8: Overview of Complete Workflow.
Appendix B NASDAQ Stock
This study selects seven stocks from the NASDAQ exchange: AAPL, AMZA, GOOGL, MSFT, NFLX, NVDA, and TSLA. These stocks represent leading companies in the technology, energy, and automotive sectors, providing high market representativeness and significant trading volumes. The NASDAQ stock dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays. This dataset reflects current market trends and serves as a timely foundation for our research. It includes daily records of opening price, closing price, highest price, lowest price, and trading volume, as well as relevant technical indicators, offering a comprehensive view of market behavior.
The training and testing periods for MACD, StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), and our system are shown in Table 7. StockFormer and TimesNet require longer training periods, with their training datasets spanning a broader time range compared to the other models. In contrast, our system does not require training and relies solely on historical stock data for making trading decisions.
Strategy Training Testing
Start End Start End
MACD 29/7/2024 30/8/2024 3/9/2024 29/10/2024
StockFormer 3/7/2023 30/8/2024 3/9/2024 29/10/2024
TimesNet 3/7/2023 30/8/2024 3/9/2024 29/10/2024
Ours 16/8/2024 30/8/2024 3/9/2024 29/10/2024
Table 7: Training and Testing Periods for Various Models on the NASDAQ Stock Dataset.
The testing period is set from September 3, 2024, to October 29, 2024, to prevent potential data leaks that could provide prior knowledge to the GPT-4o system. This timeframe ensures the fairness of the evaluation by mitigating biases from such leaks, thereby enhancing the reliability of the experimental outcomes.
Appendix C Visualization Input
Figure 7 illustrates the system’s input prompts and corresponding outputs during the strategy update process. The input prompts consist of both textual and visual components, including daily K-line charts, transaction histories, and agent trading volumes, all of which inform the strategy update.
Appendix D Simulation Process
In the Agent Trading Arena, the simulation process unfolds as follows: First, rumors are generated in the chat pool based on the previous day’s stock market analysis. Next, historical stock data is analyzed, followed by decision-making and execution. Short-term memory is formed through interactions with the environment. Finally, the system evaluates this memory, updates the strategy, and consolidates it into long-term memory. This entire process is illustrated in Figure 8.
Appendix E Simulation Process
The experiments involved several LLMs, including LLaMa-3 (Dubey et al., 2024), GPT-4o (Hurst et al., 2024), DeepSeek (Liu et al., 2024), Qwen-2.5 (Yang et al., 2024b), and Gemini-1.5 (Reid et al., 2024), corresponding to the models Meta-LLaMa-3-70B-Instruct, gpt-4o-2024-08-06, DeepSeek-chat, Qwen2.5-72B-Instruct, and Gemini-1.5-pro.