### v2

# Agent Trading Arena: A Study on Numerical Understanding in LLM-Based Agents

Abstract
Large language models (LLMs) have demonstrated remarkable capabilities in natural language tasks, yet their performance in dynamic, real-world financial environments remains underexplored. Existing approaches are limited to historical backtesting, where trading actions cannot influence market prices and agents train only on static data. To address this limitation, we present the Agent Trading Arena, a virtual zero-sum stock market in which LLM-based agents engage in competitive multi-agent trading and directly impact price dynamics. By simulating realistic bid-ask interactions, our platform enables training in scenarios that closely mirror live markets, thereby narrowing the gap between training and evaluation. Experiments reveal that LLMs struggle with numerical reasoning when given plain-text data, often overfitting to local patterns and recent values. In contrast, chart-based visualizations significantly enhance both numerical reasoning and trading performance. Furthermore, incorporating a reflection module yields additional improvements, especially with visual inputs. Evaluations on NASDAQ and CSI datasets demonstrate the superiority of our method, particularly under high volatility. All code and data are available at https://github.com/wekjsdvnm/Agent-Trading-Arena.
Agent Trading Arena: A Study on Numerical Understanding in LLM-Based Agents
Tianmi Ma1,#\#, Jiawei Du2,#\#, Wenxin Huang3, ✉, Wenjie Wang4, Liang Xie1, Xian Zhong1, ✉, and Joey Tianyi Zhou2 1 Wuhan University of Technology  2 Agency for Science, Technology and Research, Singapore 3 Hubei University  4 University of Science and Technology of China Correspondence: wenxinhuang_wh@163.com, zhongx@whut.edu.cn
footnotetext: #\# denotes equal contribution.
1 Introduction
In recent years, large language models (LLMs) have demonstrated strong capabilities in contextual understanding and generation across tasks such as natural language processing (Koshkin et al., 2024), data analysis, and reasoning (Olausson et al., 2023; Qiao et al., 2023). Their applications are rapidly extending to complex decision-making and interactive scenarios. LLM-based autonomous agents have emerged as a promising paradigm, showing notable progress in communication (Qian et al., 2024; Zhong et al., 2024), social simulation (Park et al., 2023, 2022; Lin et al., 2023), gaming (Wang et al., 2024a; Zhu et al., 2023), and multi-agent collaboration (Chen et al., 2023; Hu et al., 2023; Hong et al., 2024). These agents are increasingly regarded as a pathway toward general artificial intelligence, with growing potential in dynamic decision-making and interactive environments.
Figure 1: Comparison between Static Backtesting and the Arena. Left: Static backtesting performs one-shot analysis on historical data with fixed trading decisions, leveraging models such as Agent (market reasoning), StockFormer (predictive coding + RL hybrid), and StockMixer (time-series prediction). Right: The Arena, by contrast, supports real-time sensing, continuous adaptation, and feedback-driven execution for dynamic trading.
Despite these advances, the application of LLM-based agents in financial markets remains limited. Financial ecosystems naturally form complex multi-agent environments in which participants interact through dynamic market mechanisms. However, most existing studies (Cheng and Chin, 2024; Xiao et al., 2024) rely on historical backtesting, as illustrated in Fig.˜1, where agents cannot influence market prices and thus fail to capture interactivity and adaptive behavior. This gap underscores the need for a framework that supports real-time agent interactions and realistic market feedback.
To address these limitations, we present the Agent Trading Arena, a virtual zero-sum stock market in which LLM-based agents engage in competitive trading and directly affect price dynamics through bid-ask interactions. Our platform generates continuously evolving numerical data, forcing agents to adapt to shifting market conditions; any temporarily optimal strategy is quickly countered, rewarding adaptability and robust reasoning over memorization.
By simulating bid-ask order matching in real time, the Arena faithfully reproduces liquidity dynamics and price impact, enabling agents to experience authentic market friction and slippage. Unlike static backtests, the Arena continuously updates its state based on agents’ collective actions, creating an ever-shifting landscape that prevents overfitting to historical patterns. Because agents directly influence subsequent price trajectories, the Arena closes the loop between training and evaluation, effectively bridging the “train-test” gap of traditional financial simulations. Its modular design further supports seamless integration of new market features, such as varying transaction costs, order types, or macroeconomic shocks, making it a versatile testbed for financial AI research.
Our experiments show that LLMs struggle with numerical reasoning when processing plain-text data: they focus on absolute values, overlook percentage changes and relational patterns, and overemphasize recent trends even when earlier data are emphasized. In contrast, visual representations, such as line charts, bar graphs, and scatter plots, substantially improve both numerical reasoning and trading performance. As shown in Table˜5, chart-based inputs lead to significantly higher return rates in the Agent Trading Arena.
We further integrate a reflection module (Zhou et al., 2022; Shinn et al., 2023), which amplifies the benefits of visual data by enabling agents to iteratively refine their reasoning. This module is particularly effective with structured visual inputs, facilitating deeper pattern recognition and more strategic decision-making.
Finally, we validate our approach on real-world datasets from NASDAQ and CSI, demonstrating consistent performance gains over baselines, particularly under high-volatility market regimes.
Our contributions are summarized threefold:
  • We design the Agent Trading Arena, a dynamic and interactive simulation framework that moves beyond static backtesting with zero-sum multi-agent competition, enabling realistic evaluation of LLMs’ reasoning and adaptability.
  • We uncover LLMs’ limitations in textual numerical reasoning, highlighting difficulties with relational patterns, percentage changes, and long-range dependencies, while demonstrating superior performance with structured visual data.
  • We demonstrate that a reflection module enhances LLMs’ reasoning over complex visual inputs, leading to more accurate and strategic trading decisions.
2 Related Works
2.1 LLM-Based Agents in Dynamic Decision-Making Environments
LLMs serve as the core reasoning engines in autonomous agents, enabling planning, tool use, reflection, and adaptation within dynamic environments. Advances in NLP have empowered models such as ChatGPT (OpenAI, 2023), PaLM (Anil et al., 2023), and LLaMA (Touvron et al., 2023) to operate via natural language, making them particularly effective in interactive, text-based scenarios. Frameworks like ChatDev (Qian et al., 2024), AgentVerse (Chen et al., 2023), and MetaGPT (Hong et al., 2024) demonstrate multi-agent collaboration on software development tasks, while Voyager (Wang et al., 2024a), GROOT (Cai et al., 2024), and Optimus-1 (Li et al., 2024b) showcase autonomous exploration in open-ended environments such as Minecraft. Social simulators, including Generative Agents (Park et al., 2023) and AgentSims (Lin et al., 2023), further reveal emergent behaviors through sustained multi-agent interactions.
Together, these studies demonstrate that LLM-based agents can effectively coordinate, communicate, and make decisions in complex, evolving environments. Their adaptability and collaborative strengths highlight strong potential for real-time decision-making in finance, an information-dense and competitive domain analogous to the dynamic settings where LLMs excel.
2.2 LLMs in Financial Analysis and Decision Making
LLMs have been applied to sentiment analysis, financial question answering, and investment recommendation. Benchmarks such as SocioDojo (Cheng and Chin, 2024) evaluate financial understanding but lack interactive components. Multi-agent systems like StockAgent (Zhang et al., 2024a), TradingAgents (Xiao et al., 2024), and CryptoTrade (Li et al., 2024a) enhance performance through role specialization and iterative reflection. In contrast, traditional forecasting models, e.g., StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), PatchTST (Nie et al., 2023), PDFormer (Jiang et al., 2023), StockMixer (Fan and Shen, 2024), and iTransformer (Liu et al., 2024b), focus on price prediction using advanced methods such as wavelet transforms and temporal modeling to capture market dynamics.
However, neither paradigm addresses the “live market” paradox: without real-time adaptive mechanisms, these systems cannot simulate how participants adjust strategies in response to multi-agent interactions, news shocks, and evolving order flows. This limitation significantly reduces their practical utility in real trading environments.
Figure 2: Overall architecture of the Agent Trading Arena, consisting of three components: the Chat Pool (left) for inter-agent communication, agent modules (center) for analysis, decision-making, and reflection, and the trading platform (right), where agents’ buy/sell actions dynamically determine asset prices in a closed-loop market simulation.
3 Agent Trading Arena
To overcome the limitations of static backtesting and capture dynamic interactions among market participants, we propose the Agent Trading Arena, a closed-loop, zero-sum economic system that simulates complex quantitative scenarios in real markets (Guo et al., 2024). The overall architecture is shown in Fig.˜2 and detailed in Appendix˜A. Within this environment, agents invest in assets, earn dividends on their holdings, and pay daily expenses using virtual currency.
3.1 Virtual Market Environment
Closed-Loop System.
The Agent Trading Arena is a closed-loop virtual economy that isolates external knowledge and avoids reliance on historical data, enabling fair evaluation of agent capabilities. Unlike traditional backtesting grounded in past market records, this sandboxed system removes real-world anchors: all assets are abstract symbols, and their prices emerge solely from agent interactions through a bid-ask mechanism. By eliminating external context, the environment compels agents to adapt strategies based on internal dynamics and opponent behavior, fostering genuine decision-making and strategic evolution.
Bid-Ask Mechanism.
Asset prices arise exclusively from a bid-ask process: each agent’s orders directly influence the market price, with no external benchmarks. This closed-loop design rigorously captures endogenous price formation, faithfully modeling liquidity constraints, order book dynamics, market depth, and slippage. Every execution reshapes the market microstructure, producing non-trivial price impacts and path-dependent outcomes that evolve with aggregate agent behavior over time.
3.2 Chat Pool
To simulate real-world information flow, we introduce a shared Chat Pool where agents post and read “news” items, analyst commentary, and social media-style opinions. By reproducing both informative signals and noisy interference, including hallucinations (Gekhman et al., 2024) and misinformation, the Chat Pool reflects how public information can shape, and at times distort, traders’ narratives and expectations. This process increases the complexity of strategic reasoning and adaptive behavior in the market.
3.3 Market Mechanisms
We incorporate two key incentives to sustain active trading and realistic dynamics: 1) Dividends & Capital Gains: Agents earn income from both price appreciation and dividends, distributed proportionally to holdings. This implicitly anchors prices and rewards low-cost positions. 2) Daily Wealth Fee: A cost proportional to each agent’s total wealth is charged daily. Only those earning sufficient dividends can cover this fee, which discourages passive holding and promotes frequent, strategic trades.
Together, these mechanisms maintain continuous market activity, encourage adaptive decision-making, and prevent reliance on static or exploitative strategies in a zero-sum environment.
3.4 Zero-Sum Competition
By design, the Arena is zero-sum: the profit of one agent is the loss of another, meaning there is no fixed “best” strategy. Competitors counter any temporarily successful tactic, compelling agents to learn implicit rules, interpret delayed numerical feedback, and continuously adapt in real time. With results unfolding gradually and misinformation potentially introduced by rivals, agents must rely on experiential learning and develop context-aware strategies through ongoing adaptation rather than predefined rules or external knowledge.
4 Proposed Method
Our agent, ArenaTrader, integrates textual and visual modalities to form a comprehensive understanding of market conditions. It employs a structured three-stage framework, market analysis, trade execution, and strategic reflection, to leverage this multi-modal representation. This design enables informed and adaptive decision-making based on data-driven insights and iterative feedback.
Figure 3: Textual vs. Visual Input–Output Representations. Left: Agent buy/sell records, daily trade prices, and corresponding candlestick charts. Top right: Visual-input outputs highlight overall trends and long-term patterns. Bottom right: Textual-input outputs emphasize specific price points.
4.1 Numerical Data Representations
Textual Input.
Time-series stock data are represented as plain text, including timestamps, prices, and volumes, to provide explicit semantics (Hu et al., 2024; Zhang et al., 2024b). However, LLMs often fixate on absolute values, overlook percentage changes, and overemphasize recent entries, which limits their ability to reason across long horizons (see lower right of Fig.˜3).
Visual Input.
To complement textual data, we introduce visualizations such as line charts, bar graphs, and scatter plots (Yang et al., 2025), which highlight global trends, relational patterns, and volatility at a glance (see upper right of Fig.˜3). These visualizations are consistent with the textual input, ensuring a cohesive representation. This structured format enables LLMs to integrate local details with broader market context more effectively.
4.2 Agent Architecture
Investment Analysis.
Inspired by human trading behavior, each agent begins by analyzing market conditions before executing trades. This process involves extracting and synthesizing multi-modal signals from the Chat Pool, price histories, and fundamental indicators. By combining these heterogeneous inputs, the agent constructs a stock ranking framework based on predefined heuristics or learned criteria, allowing it to identify promising assets for portfolio allocation.
Action Decision.
Agents execute multiple trades per day, making buy, sell, or hold decisions by integrating diverse signals, including analysis reports, real-time market data, and historical performance trends. They evaluate volatility, price momentum, and portfolio risk exposure to compute optimal trade volumes under capital constraints. Orders are executed via the bid-ask mechanism, with portfolios updated accordingly. Each decision aims to maximize long-term returns while maintaining a competitive edge in the evolving virtual market.
Figure 4: Design of the Reflection Module. Left: Agent trading records, strategy evaluations, and top- and bottom-performing strategies from the strategy library. Top right: Corresponding stock chart. Bottom right: Newly generated strategy based on current information.
Strategy Optimization via Reflection.
At the end of each trading day, the reflection module distills daily experiences into bidirectional learning signals (see Fig.˜4). It first evaluates the effectiveness of trading trajectories and strategies using a scoring function. Then, by contrasting top- and bottom-performing strategies, agents generate refined action plans and archive unsuccessful tactics in a strategy library. This iterative process drives continuous strategy evolution and more robust decision-making.
5 Experimental Results
5.1 Experimental Setup
Environment.
The Agent Trading Arena is lightweight and accessible, requiring only API credits and no GPU for deployment. Details on the selection of different LLMs are provided in Appendix˜D.
ID Name
Duration
Initial
Capital
Profession
1 Amy 1 100,000 AI Researcher
2 Bruce 2 100,000 Lawyer
3 Charles 1 100,000 Doctor
4 David 3 100,000 Engineer
5 Ella 2 100,000 Teacher
6 Frank 5 100,000 Entrepreneur
7 Grace 4 100,000 Accountant
8 Hank 2 100,000 Architect
9 Ivy 3 100,000 Marketing Manager
Table 1: Agent Characteristics. Duration denotes each agent’s investment horizon, and Initial Capital indicates the starting funds for trading.
ID Ticker DPS Closing Prices Initial Quantity
1 A 22 454.17, …, 445.60 1,200
2 B 23 354.17, …, 465.80 1,000
3 C 25 500.47, …, 440.60 1,600
Table 2: Stock Details. DPS denotes dividend per share; Closing Prices are historical end-of-day closing prices; and Initial Quantity specifies the starting number of shares.
Datasets.
We configure the Arena with varying numbers of agents and stocks, all initialized with the same capital (see Tables˜1 and 2). To validate our agents, we simulate portfolio investments on subsets of NASDAQ and CSI datasets. For NASDAQ, we run a two-month simulation (September 3–October 29, 2024) with an initial capital of 100,000 units, excluding weekends and holidays. For CSI, we conduct a one-year simulation (January 2–December 31, 2024) with the same capital, also excluding non-trading days. Both datasets, collected from Yahoo Finance and Baostock, provide daily records of opening, closing, high, and low prices, trading volume, and associated technical indicators, offering a holistic and up-to-date depiction of market behavior. Additional details are given in Appendix˜B.
Baselines.
We compare ArenaTrader against the following strategies:
1) Buy & Hold: A passive strategy that purchases assets initially and holds them throughout the period without further trading.
2) SMA (Gencay, 1996): A moving average crossover strategy used to detect bullish and bearish momentum. We evaluate window sizes [5, 10, 15, 20] and select the best based on validation results.
3) ZMR (Poterba and Summers, 1988): A zone-based reversal strategy that trades on deviations from preset bounds, with thresholds and holding periods tuned via validation.
4) MACD (Chong and Ng, 2008): A momentum strategy based on MACD and signal line crossovers, where MACD is the 12-day EMA minus the 26-day EMA, and the signal line is the 9-day EMA.
5) StockFormer (Gao et al., 2023): A hybrid model combining predictive coding and reinforcement learning to capture future dynamics and asset correlations. We follow the default setup from the original implementation.
6) TimesNet (Wu et al., 2023): Transforms 1D time series into 2D representations to model multi-period temporal variations. We follow the original setup and select the look-back window from [5, 10, 15, 20] based on validation.
7) StockMixer (Fan and Shen, 2024): A lightweight MLP that mixes indicators, time, and stocks for efficient modeling of market dynamics. We follow the settings specified in the original work.
Metrics.
We evaluate using five metrics:
1) Total Return (TR): $(C_{1}-C_{0})/C_{0}$, where $C_{0}$ and $C_{1}$ are initial and final capital.
2) Win Rate (WR): $N_{w}/N_{t}$, the ratio of profitable trading days to total trading days.
3) Sharpe Ratio (SR): $(R_{p}-R_{f})/\sigma_{p}$, where $R_{p}$ is the mean daily return, $\sigma_{p}$ is the standard deviation of daily returns, and $R_{f}$ is set to 0 to emphasize risk-adjusted returns.
4) Mean Daily Return (Mean): Average daily return over the simulation.
5) Return Volatility (Std): Standard deviation of daily returns, measuring risk.
5.2 Experimental Results
In the Agent Trading Arena, we configure multiple agents and LLM backbones to promote diversity. The primary agent, ArenaTrader, consistently uses our full-featured design, including visual perception and reflection, enabling focused analysis. We evaluate ArenaTrader on real-world datasets, with corresponding prompts detailed in Appendix˜C.
Model Avg. Trend TR (%) $\uparrow$ Mean (%) $\uparrow$ Std (%) $\downarrow$ WR (%) $\uparrow$ SR $\uparrow$ $\Delta$ (%) $\uparrow$
ArenaTrader w/ Qwen-VL-128k -0.4632 33.3018 2.6526 1.0150 100.00 2.6135 33.7650
ArenaTrader w/ GPT-4o-mini 5.4722 38.6460 3.0249 1.4235 100.00 2.1249 33.1738
ArenaTrader w/ Gemini-1.5 7.2743 41.3264 3.1946 0.1588 100.00 20.1128 34.0521
ArenaTrader w/ Qwen-VL-32k -0.5216 35.8150 3.1213 1.6079 90.00 1.9412 36.3366
ArenaTrader w/ GPT-4o 3.5881 47.6851 3.6095 0.5327 100.00 6.7765 44.0970
Table 3: Investment Performance of ArenaTrader with Different LLMs. Avg. Trend denotes the average market trend, and $\Delta$ indicates the gain relative to this trend. GPT-4o achieves the highest returns and stability, demonstrating superior data-driven reasoning among vision-capable LLMs. Bold and underlined values denote the best and second-best results, respectively.
Refer to caption Refer to caption
(a) Gemini-1.5 (b) Qwen-VL-128k
Refer to caption Refer to caption
(c) GPT-4o-mini (d) GPT-4o
Figure 5: ArenaTrader Performance vs. Market Volatility. Total return of ArenaTrader in the Agent Trading Arena, compared with the most volatile single stock and the average return across all stocks. This comparison highlights the agent’s robustness under varying market conditions.
Vision-Enhanced Agent Performance.
Tables˜3 and 5 report the returns of ArenaTrader under different vision-capable LLMs. GPT-4o (Hurst et al., 2024) achieves the highest TR, followed by Gemini-1.5 (Reid et al., 2024). For volatility-adjusted returns, GPT-4o again leads, with Qwen-VL-32k (Wang et al., 2024b) second. These results highlight GPT-4o’s strong data-driven reasoning and robustness across varying volatility regimes, underscoring its ability to integrate visual and numerical signals in dynamic markets.
Strategy NASDAQ CSI
TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$ TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
Buy & Hold 6.06 0.155 0.916 61.54 0.169 0.73 0.004 0.403 43.15 0.010
SMA 0.19 0.006 0.561 46.15 0.011 -3.45 -0.014 0.319 33.20 -0.044
ZMR 1.04 0.030 0.814 41.03 0.037 1.24 0.006 0.361 41.91 0.016
MACD 5.62 0.146 1.033 58.97 0.141 -0.67 -0.002 0.373 40.66 -0.006
StockFormer (Gao et al., 2023) 11.23 0.281 0.121 56.41 0.231 14.83 0.636 1.128 41.08 0.056
TimesNet (Wu et al., 2023) 11.41 0.283 1.035 56.41 0.273 32.93 0.139 2.053 48.55 0.068
StockMixer (Fan and Shen, 2024) 14.64 0.357 1.173 56.41 0.305 43.00 0.153 0.907 27.80 0.168
ArenaTrader w/ GPT-4o (Ours) 15.99 0.450 1.110 58.97 0.348 48.93 0.176 1.432 51.45 0.123
Table 4: Strategy Performance on NASDAQ and CSI. Our method demonstrates strong robustness and generalization, achieving near-optimal performance without additional training.
Evaluation on Real-World Data.
Using GPT-4o as the backbone, Table˜4 presents results on NASDAQ and CSI. Training details for StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), and StockMixer (Fan and Shen, 2024) are provided in Appendix˜B. Despite extensive training, these models underperform our method, which uses a 10-day input window selected via ablation. The Sharpe Ratio (SR) of NASDAQ-100 benchmark during the same period is 0.189, while CSI 300 records 0.054. In contrast, our model achieves 0.348 on NASDAQ-100 (an improvement of 0.159) and 0.123 on CSI 300 (an improvement of 0.069).
(a) NASDAQ
(b) CSI
Figure 6: Backtesting Results on Real-World Stock Datasets. Two-month backtest on NASDAQ and one-year backtest on CSI. The proposed method consistently delivers near-optimal performance and outperforms all baselines, particularly during periods of high market volatility.
Building on these results, Fig.˜6 provides a detailed analysis of portfolio trajectories and market conditions, illustrating our agent’s near-optimal performance, especially during high-volatility periods. Notably, our approach operates without task-specific retraining, relying solely on historical price data. This design underscores the robustness and strong generalization ability of our agent in real-world, dynamic market environments.
5.3 Ablation Studies
We conduct ablation studies in the Arena to evaluate the impact of input formats and the reflection mechanism on agent performance. To further assess the effect of sequence length across textual and visual modalities, we perform controlled backtesting on NASDAQ. Backtesting mitigates dynamic data fluctuations, allowing sequence length to be isolated as the primary variable.
Model Textual Visual TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
w/o Reflection
LLaMa-3 (Dubey et al., 2024) 4.09 0.686 2.270 62.50 0.196
DeepSeek (Liu et al., 2024a) 27.31 2.248 2.755 92.21 1.530
Qwen-2.5 (Yang et al., 2024) 30.37 2.680 1.459 93.75 1.703
Gemini-1.5 (Reid et al., 2024) 14.32 1.221 2.778 83.84 1.729
19.04 1.604 1.165 90.91 1.376
23.96 1.981 1.334 100.00 1.485
GPT-4o (Hurst et al., 2024) 13.04 1.139 1.929 54.55 0.590
17.07 1.454 1.504 72.73 0.967
26.18 2.157 2.058 90.91 1.048
w/ Reflection
LLaMa-3 (Dubey et al., 2024) 10.25 1.667 2.403 66.67 0.694
DeepSeek (Liu et al., 2024a) 30.62 2.470 1.512 100.00 1.633
Qwen-2.5 (Yang et al., 2024) 38.91 3.324 1.073 100.00 3.099
Gemini-1.5 (Reid et al., 2024) 29.45 2.519 3.012 93.51 2.210
37.01 2.911 1.235 100.00 2.357
41.33 3.195 0.159 100.00 20.113
GPT-4o (Hurst et al., 2024) 33.65 2.671 2.171 98.70 2.142
35.76 2.821 0.678 100.00 4.159
47.69 3.610 0.533 100.00 6.777
Table 5: Performance Comparison with and without Reflection Using Textual and Visual Inputs. Agents using visual data consistently outperform those using text alone, while combining both modalities yields the strongest results.
Textual vs. Visual Inputs.
We compare agents using textual-only inputs with those leveraging visualizations (scatter plots, line charts, bar graphs). Non-vision LLMs receive text only without reflection, while vision-capable LLMs are evaluated under text-only, visual-only, and combined-input settings. Corresponding prompts are provided in Appendix˜C, and details of LLM choices are in Appendix˜D. Results in Table˜5 show that visual inputs substantially improve returns, with combined inputs achieving the best overall performance.
Impact of Reflection.
Incorporating the reflection module further enhances performance across all LLMs. For GPT-4o, TR improves from 33.65% with text-only to 47.70% with visual input, a relative gain of 41.7%. Beyond TR, reflection also boosts SR and WR, indicating higher profitability, stability, and consistency in trading outcomes. These findings confirm that strategic reflection, particularly with visual data, strengthens agents’ reasoning and decision-making under volatile market conditions.
Input Window TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
Textual 5 days 3.63 0.102 1.588 62.5 0.064
10 days 5.58 0.145 1.334 60.0 0.109
15 days 6.47 0.171 1.681 57.5 0.102
20 days 1.67 0.050 1.296 62.5 0.039
Visual 5 days 8.21 0.207 1.396 57.5 0.149
10 days 6.90 0.174 1.164 67.5 0.149
15 days 6.90 0.174 1.216 60.0 0.143
20 days 2.81 0.075 1.101 62.5 0.068
Textual
Visual
5 days 14.33 0.372 1.719 60.0 0.216
10 days 15.99 0.450 1.110 59.0 0.348
15 days 14.73 0.380 1.684 57.5 0.226
20 days 13.51 0.352 1.577 65.0 0.223
Table 6: Performance with Textual and Visual Inputs Across Different Window Sizes. Window denotes the number of consecutive trading days provided as historical input before the current trading day. Textual + Visual consistently achieves the best performance, particularly with a 10-day window, highlighting the importance of both modality choice and time-window design.
Impact of Time-Series Length.
Experiments with GPT-4o as the backbone are summarized in Table˜6, which reports performance across modalities and time windows. Overall, “Textual + Visual” consistently outperforms single modalities, particularly with a 10-day window, achieving the highest TR (15.99%), SR (0.348), and Mean (0.450%). Visual inputs perform best in the 5–10 day range, offering strong returns, stability, and win rates. Performance declines at 20 days, likely due to noise accumulation or overfitting. These results underscore the importance of both multi-modal fusion and careful time-window selection for robust time-series decision-making.
6 Conclusion
We presented the Agent Trading Arena, a closed-loop, zero-sum simulation framework designed to model complex market dynamics and evaluate LLMs on numerical reasoning tasks. Our experiments show that while LLMs struggle with plain-text numerical data, they achieve significantly higher returns when supported by visual representations, underscoring the importance of charts and graphs in complex decision-making. Integrating a reflection module further improves performance by enabling iterative analysis and strategic refinement. Validation on both NASDAQ and CSI confirms that vision-enhanced LLM agents consistently outperform text-only counterparts without additional training. Overall, this work highlights both the capabilities and limitations of LLMs in numerical reasoning and establishes the Agent Trading Arena as a robust testbed for advancing research in adaptive, real-world financial applications.
Limitations
This study investigates the visual reasoning abilities of LLMs in a virtual stock trading environment. While such a controlled setup enables systematic analysis, it inevitably simplifies many aspects of real-world financial systems, including uncertainty, feedback loops, and human interactions. Bridging the gap between virtual environments and real-world applications therefore remains a significant challenge.
A promising direction for future work is to explore how agents can be gradually integrated into real financial workflows, interacting with live data, adapting to real-time signals, and contributing meaningful decisions or strategies. This would move beyond passive evaluation toward active participation, where agents not only respond to environments but also generate novel behavioral data and insights to inform downstream analysis, model training, and system design. Expanding the task scope and incorporating more diverse modalities may further support the development of generalizable and robust evaluation frameworks.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62301213 and 62271361, the Hubei Provincial Key Research and Development Program under Grant 2024BAB039, and the Open Project Funding of the Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University, under Grant 2024BDIAA01.
Appendix A Agent Trading Arena
A.1 Overview
As illustrated in Fig.˜7, each agent’s workflow integrates LLMs for chat pool interactions, stock analysis, decision-making, and reflection. In the stock analysis and decision-making modules, all outputs are validated for consistency with both common sense and operational requirements before execution.
Figure 7: Agent Workflow Components: Share Trading involves stock analysis, decision validation, and trade execution; Environment and Memory manage memory and process trade orders; and Reflection focuses on strategy assessment and refinement based on feedback.
A.1.1 Action Decision-Making
Action generation follows the LLM framework. The agent responsible for generating actions receives corresponding prompts via SQLite. Based on these prompts and specified output formats, the agent decides whether to buy, sell, or hold stocks. The action generation process is outlined in Eq.˜1, with input prompts shown in Fig.˜8 and the corresponding outputs displayed in the adjacent figure.
$\begin{cases}A_{\mathrm{date}+1}^{t}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{t},S_{\mathrm{date}}\right),&\text{if~}t\text{~is Iters},\\ A_{\mathrm{date}}^{t+1}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{t},S_{\mathrm{date}}\right),&\text{otherwise},\end{cases}$ (1)
where $\mathrm{Ins}$ represents the environment introduction, $Z_{\mathrm{date}}^{t}$ denotes the memory of the stock transaction on day $\mathrm{date}$ retrieved from the database, and $S_{\mathrm{date}}$ is the strategy for day $\mathrm{date}$ generated via reflection.
Figure 8: Inputs and Outputs in Action Decision-Making.
A.1.2 Environmental Interaction
To isolate external influences, we created a virtual sandbox environment where each agent is assigned a unique ID, and their actions affect the environment. The function $\phi$ facilitates environmental interactions, as shown in Algorithm˜1, where “OPS” retrieves agent actions, “date” refers to the trading date, and “Z” represents the memory used for interaction with the environment. Through $\phi$, each agent’s actions, such as buying or selling stocks, determine the current stock price and update the trading platform, including stock prices and available shares. Stock prices are independent of external factors and are influenced solely by the sandbox’s internal dynamics. The stock price is updated with each transaction according to the following formula:
$\mathrm{Price}_{\mathrm{curr}}$ $=\delta\left(Q,F,\mathrm{Price}_{\mathrm{curr}},\mathrm{Price}_{\mathrm{deal}}\right)$ (2)
$=\frac{\mathrm{Price}_{\mathrm{deal}}\cdot Q\cdot F+\mathrm{Price}_{\mathrm{curr}}\cdot Q_{\mathrm{total}}}{Q\cdot F+Q_{\mathrm{total}}},$
where $Q$ is the quantity of stock traded, $F$ is the fluctuation constant, $\mathrm{Price}_{\mathrm{curr}}$ is the current stock price, $Q_{\mathrm{total}}$ is the total number of shares available, and $\mathrm{Price}_{\mathrm{deal}}$ is the price at which the trade occurs.
Input: OPS: Function to retrieve agents’ actions, date: The current trading date, Z: Memory used for interaction with the environment
Output: Z
for $P\in\text{Persons}$ do
 $A\leftarrow\text{OPS}(t,P)$;
 $O,N,Q,\text{Price}_{\text{deal}}\leftarrow\text{Extract}(A)$;
 $\text{Price}_{\text{curr}},Q_{\text{total}}\leftarrow\text{Stocks}(N)$;
 $\text{Price}_{\text{curr}}\leftarrow\delta(Q,F,\text{Price}_{\text{curr}},\text{Price}_{\text{deal}})$;
 if $O=$ “buy” then
    $\text{Cash}\leftarrow\text{Price}_{\text{curr}}\cdot Q$;
    if $\text{Cash}<P.\text{Cash}$ then
       $Z\leftarrow\text{SubmitOrder}(O,N,t,\text{Price}_{\text{curr}},Q)$;
 if $O=$ “sell” then
    $\text{Hold}\leftarrow P(N)$;
    if $\text{Hold}\neq\text{None}$ then
       $Q_{N}\leftarrow\text{Hold}[\text{``Q''}]-Q$;
       if $Q_{N}>0$ then
          $Z\leftarrow\text{SubmitOrder}(O,N,t,\text{Price}_{\text{curr}},Q)$;
$\text{Market}(\text{date},\text{Persons})$;
return Z;
Algorithm 1 Environmental Interaction
The function $\phi$ executes trading orders and updates the stock price in real-time based on the agent’s actions. To prevent excessive volatility and mitigate risk during trading cycles, a daily price fluctuation cap is enforced. Before executing any transaction, each agent evaluates its available funds and refrains from proceeding if insufficient capital is available.
A.1.3 Memory
The superior performance of LLM-based agents arises from the extensive internal knowledge acquired during pre-training. The large number of parameters in LLMs enables the retrieval of diverse information and supports logical and inferential reasoning. To further enhance knowledge retrieval across various tasks, we incorporate a memory module that empowers LLM-based agents with self-improvement capabilities. This memory module facilitates strategy reflection through time-series feedback. Unlike qualitative tasks, quantitative feedback evolves incrementally with subtle differences, presenting a challenge for the generalization of existing LLM-based agents.
To minimize the influence of pre-existing knowledge, we assign specific roles to each agent. The memory module records sensory inputs and accumulates valuable experiences based on immediate feedback following actions. These experiences are stored in a database for future reference. The interaction history between an agent’s behavior and the environment constitutes short-term memory, enabling the agent to retain recent events. The historical trajectory is defined as:
$M_{\mathrm{date}}^{t}=\zeta\left(I_{\mathrm{date}}^{t},\mathrm{Out}_{\mathrm{date}}^{t}\right),$ (3)
where $\zeta$ processes key information, $I_{\mathrm{date}}^{t}$ represents the input prompt, and $\mathrm{Out}_{\mathrm{date}}^{t}$ denotes the resulting output.
The day’s trajectory constitutes short-term memory, which is expressed as:
$Z_{\mathrm{date}}^{t}=\left(M_{\mathrm{date}}^{0},M_{\mathrm{date}}^{1},\dots,M_{\mathrm{date}}^{t}\right).$ (4)
The process for updating short-term memory is given by:
$Z_{\mathrm{date}}^{t+1}=M_{\mathrm{date}}^{t}\cup Z_{\mathrm{date}}^{t},\quad t\in\{0,1,\dots,T\},$ (5)
where $Z_{\mathrm{date}}^{t}$ represents the short-term memory, and $T$ denotes the maximum number of iterations.
The reflection model serves as long-term memory, enabling self-reflection and the consolidation of knowledge.
A.1.4 Reflection
We propose a strategy distillation method that transforms quantitative results into descriptive text, which is then used as prompts for LLMs. This approach aids in the analysis of results and the generation of actionable, qualitative summaries, enabling LLMs to derive new strategies. These strategies are implemented, monitored, and evaluated over time, while underperforming strategies are archived for future review.
Initially, we evaluate the day’s trajectory memory and associated strategies. The evaluation function is defined as:
$E_{\mathrm{date}}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{T},S_{\mathrm{date}}\right),$ (6)
where $\Psi$ represents the evaluation function based on LLMs, $\mathrm{Ins}$ contains the agent role descriptions and output requirements, and $Z_{\mathrm{date}}^{T}$ denotes the memory for the given day.
Based on the evaluation results and previous strategies, we generate the latest strategy. Past strategies are stored in a library and scored. For the new strategy, we select the top five best-performing and the bottom five worst-performing strategies to provide both positive and negative feedback. The strategy update formula is:
$S_{\mathrm{date}+1}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{T},E_{\mathrm{date}},S_{\mathrm{date}}\right),$ (7)
where $E_{\mathrm{date}}$ represents the evaluation results for the day, and $S_{\mathrm{date}}$ is the strategy for that day.
Long-term memory is generated through reflection, as represented by:
$Z^{\prime}_{\mathrm{date}}=\left(S_{0},S_{1},\dots,S_{\mathrm{date}}\right).$ (8)
The process for updating long-term memory is defined as follows:
$Z^{\prime}_{\mathrm{date}+1}=S_{\mathrm{date}}\cup Z^{\prime}_{\mathrm{date}},$ (9)
$\mathrm{date}\in\{0,1,\dots,\mathrm{DAYS}\}.$
where $Z^{\prime}_{\mathrm{date}}$ represents the long-term memory, and $\mathrm{DAYS}$ denotes the maximum number of days.
Together, short-term and long-term memory provides essential context for the agents. Success in this environment depends on their ability to understand the game rules and develop strategies that outmaneuver competitors. Agents continuously refine their strategies based on incremental quantitative feedback, adjusting their actions to align with long-term objectives.
Appendix B NASDAQ and CSI
This study selects seven stocks from the NASDAQ exchange: AAPL, AMZA, GOOGL, MSFT, NFLX, NVDA, and TSLA. These stocks represent leading companies in the technology, energy, and automotive sectors, providing high market representativeness and significant trading volumes. The NASDAQ dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays. This dataset reflects current market trends and serves as a timely foundation for our research. It includes daily records of opening price, closing price, highest price, lowest price, and trading volume, as well as relevant technical indicators, offering a comprehensive view of market behavior.
This study also selects seven representative stocks from the Chinese stock market (CSI), covering various sectors with industry-leading companies: Kweichow Moutai (600519.SH), China Merchants Bank (600036.SH), CATL (300750.SZ), BYD (002594.SZ), LONGi Green Energy (601012.SH), ZTE (000063.SZ) and Mindray (300760.SZ). These stocks span consumer goods, finance, new energy, telecommunications, and healthcare, representing leading companies with strong market influence, stable performance, and innovative potential. For CSI, we conducted a one-year simulation from January 2 to December 31, 2024, using the same initial capital setup, while excluding weekends and official trading holidays. The data set is sourced from Baostock and includes daily records such as the opening price, closing price, highest price, lowest price, trading volume, and multiple technical indicators. This comprehensive data set provides a solid foundation for analyzing market trends and investment strategies within the CSI framework.
The training and testing periods for StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), StockMixer (Fan and Shen, 2024), and our system is shown in Table˜7. StockFormer, TimesNet, and StockMixer require longer training periods, and their training data sets span a broader time range compared to the other models. In contrast, our system does not require training and relies solely on historical stock data to make trading decisions.
Strategy NASDAQ CSI
Training Testing Training Testing
Start End Start End Start End Start End
Buy & Hold w/o 25/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 21/12/2023 29/12/2023 2/1/2024 31/12/2024
SMA w/o 1/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 1/12/2023 29/12/2023 2/1/2024 31/12/2024
ZMR w/o 10/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 8/12/2023 29/12/2023 2/1/2024 31/12/2024
MACD w/o 4/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 4/12/2023 29/12/2023 2/1/2024 31/12/2024
StockFormer w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
TimesNet w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
StockMixer w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
Ours w/o 16/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 1/12/2023 29/12/2023 2/1/2024 31/12/2024
Table 7: Training and Testing Periods for Various Models on NASDAQ and CSI.
The testing period is set from September 3, 2024, to October 29, 2024, to prevent potential data leaks that could provide prior knowledge to the GPT-4o system. This timeframe ensures the fairness of the evaluation by mitigating biases from such leaks, thereby enhancing the reliability of the experimental outcomes.
We also select two stocks from the CSI dataset, each representing a different growth pattern, for simulation. As shown in Table˜8, our method consistently outperforms the average trend of the selected stocks.
Trend LLMs TR $\uparrow$ SR $\uparrow$
Average Min Max
-5.86 -13.77 -0.21 GPT4o -0.50 -0.021
Qwen 2.01 0.207
LlaMa3 1.44 0.122
11.37 -4.25 26.05 GPT4o 20.61 0.465
Qwen 16.46 0.420
LlaMa3 4.14 0.167
Table 8: Performance on CSI Stocks with Different Growth Patterns. Our method consistently outperforms the average trend of both selected stocks.
Appendix C Prompt Design Across Modalities
The prompts used for stock analysis with text-only, vision-only, and combined inputs are shown below. Additional prompts for others can be found at: https://github.com/wekjsdvnm/Agent-Trading-Arena.
Text-only Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge. - You will be given information related to specific stocks, the whole market, and your existing investments, which is used to analyze the current stock market and investment situation. - Your overall objective is to make as much profit as possible. Stock information: - A: - The closing prices in the past 10 days are: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - Dividend per share: 22 - Current price change: +0.00%, Current price: 511.79 - Intraday High: 511.79 - Intraday Low: 511.79 - Intraday Mean: 511.79 - B: - The closing prices in the past 10 days are: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - Dividend per share: 23 - Current price change: +0.00%, Current price: 480.61 - Intraday High: 480.61 - Intraday Low: 480.61 - Intraday Mean: 480.61 - C: - The closing prices in the past 10 days are: [455.90, 440.532, 424.91, 419.75, 420.48, 421.31, 420.12, 421.09, 435.33, 435.33] - Dividend per share: 25 - Current price change: +-0.00%, Current price: 435.33 - Intraday High: 435.33 - Intraday Low: 435.33 - Intraday Mean: 435.33 Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B. This merger is rumored to lead to significant synergies, possibly resulting in a stock price surge for both A and B. Consequently, many traders anticipate a rise in demand for these stocks, suggesting a timely investment could yield remarkable returns. Meanwhile, analysts rave about Stock C’s steady position and attractive dividends, marking it as a fortress stock that continues to deliver…… Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - B: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - The prices in the past 10 days are: prices: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - current price change: +0.00%, current price: 480.61, cost price: 455.8 - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT - The prices in the past 10 days are: prices: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - current price change: +0.00%, current price: 511.79, cost price: 468.98 Investment strategy: try to maximize profit. Task: - Based on (but not limited to) the following information, you need to summarize and conclude the 3 most important and valuable analysis results from the above information. The analysis results must fully fit the focus of the investment strategy, and each result should describe the relationship between its content and investment strategy. In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) Output the response to the prompt above in JSON. Each analysis result should be started with "-", and ended with a line break. Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
Visual-only Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge. - You will be given information related to specific stocks, the whole market, and your existing investments, which is used to analyze the current stock market and investment situation. - Your overall objective is to make as much profit as possible. Stock information: [Uncaptioned image] (a) Stock A Price [Uncaptioned image] (b) Stock B Price [Uncaptioned image] (c) Stock C Price [Uncaptioned image] (d) Buy & Sell Assert [Uncaptioned image] (e) Trading Record Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B. This merger is rumored to lead to significant synergies, possibly resulting in a stock price surge for both A and B. Consequently, many traders anticipate a rise in demand for these stocks, suggesting a timely investment could yield remarkable returns. Meanwhile, analysts rave about Stock C’s steady position and attractive dividends, marking it as a fortress stock that continues to deliver consistent profits with minimal risk. As traders weigh their options, strategic portfolio rebalancing tips towards these dividends-rich stocks seem to be the talk of the town. However, with the volatile nature of stock rumors, it remains essential to tread carefully and keep an eye on reliable sources for confirmation. Will this buzz hold substance, or is it merely hot air? Only time will reveal the truth behind the whispers, leaving the market in suspense. Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - C: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT Investment strategy: try to maximize profit. Task: - Based on (but not limited to) the following information, you need to summarize and conclude the 3 most important and valuable analysis results from the above information. The analysis results must fully fit the focus of the investment strategy, and each result should describe the relationship between its content and investment strategy. In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) - The provided stock information visualizations. Output the response to the prompt above in JSON. Each analysis result should be started with "-", and ended with a line break. Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
Combined-Input Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge. - You will be given information related to specific stocks, the whole market, and your existing…… Stock information: - A: - The closing prices in the past 10 days are: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - Dividend per share: 22 - Current price change: +0.00%, Current price: 511.79 - Intraday High: 511.79 - Intraday Low: 511.79 - Intraday Mean: 511.79 - B: - The closing prices in the past 10 days are: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - Dividend per share: 23 - Current price change: +0.00%, Current price: 480.61 - Intraday High: 480.61 - Intraday Low: 480.61 - Intraday Mean: 480.61 - C: - The closing prices in the past 10 days are: [455.90, 440.532, 424.91, 419.75, 420.48, 421.31, 420.12, 421.09, 435.33, 435.33] - Dividend per share: 25 - Current price change: +-0.00%, Current price: 435.33 - Intraday High: 435.33 - Intraday Low: 435.33 - Intraday Mean: 435.33 [Uncaptioned image] (a) Stock A Price [Uncaptioned image] (b) Stock B Price [Uncaptioned image] (c) Stock C Price [Uncaptioned image] (d) Buy & Sell Assert [Uncaptioned image] (e) Trading Record Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B. This merger is rumored to lead to significant synergies…… Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - B: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - The prices in the past 10 days are: prices: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - current price change: +0.00%, current price: 480.61, cost price: 455.8 - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT - The prices in the past 10 days are: prices: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - current price change: +0.00%, current price: 511.79, cost price: 468.98 Investment strategy: try to maximize profit. Task: - Based on (but not limited to) the following information, …… describe the relationship between its content and investment strategy. In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) - The provided stock information visualizations. Output the response to the prompt above in JSON. Each analysis result should be started with "-",…… Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
Appendix D Simulation Process
The experiments involved several LLMs, including LLaMa-3 (Dubey et al., 2024), GPT-4o (Hurst et al., 2024), GPT-4o-mini (Hurst et al., 2024), DeepSeek (Liu et al., 2024a), Qwen-2.5 (Yang et al., 2024), Qwen-VL-32k (Wang et al., 2024b), Qwen-VL-128kk (Wang et al., 2024b), and Gemini-1.5 (Reid et al., 2024), corresponding to the models Meta-LLaMa-3-70B-Instruct, gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, DeepSeek-chat, Qwen2.5-72B-Instruct, qwen-vl-max, qwen-vl-max-latest, and Gemini-1.5-pro.
Appendix E Visualization Input
Fig.˜9 illustrates the system’s input prompts and corresponding outputs during the strategy update process. The input prompts consist of both textual and visual components, including daily K-line charts, transaction histories, and agent trading volumes, all of which inform the strategy update.
Figure 9: Visualization of Stock Inputs and Corresponding Trading Strategy Outputs.
Appendix F Simulation Process
In the Agent Trading Arena, the simulation process unfolds as follows: First, rumors are generated in the chat pool based on the previous day’s stock market analysis. Next, historical stock data is analyzed, followed by decision-making and execution. Short-term memory is formed through interactions with the environment. Finally, the system evaluates this memory, updates the strategy, and consolidates it into long-term memory. This entire process is illustrated in Fig.˜10.
Figure 10: Overview of Complete Workflow.
Impact of Modality on LLM Competitiveness.
We employed a relative evaluation method for this experiment. The first and second agents used various LLMs in textual and visual settings, respectively, while the remaining agents were based on LLaMa-3 (Dubey et al., 2024) as the baseline. This setup aimed to explore the impact of different agents and modalities on LLM performance. The results are shown in Table˜9. The findings indicate that DeepSeek (Liu et al., 2024a) exhibited stronger competitive performance across different LLM environments, suggesting unique strengths or optimizations that enable it to adapt more effectively to the task’s complexities.
LLMs T V TR $\uparrow$ SR $\uparrow$
LLaMa-3 100.00 100.00
DeepSeek + 35.03 + 45.63
Gemini-1.5 + 10.99 + 14.93
+ 11.49 + 29.47
GPT-4o + 10.39 + 28.29
+ 17.18 + 40.77
Table 9: Performance Comparison in Trading Decisions w/o Reflection, Competing Pairwise with LLaMa-3. DeepSeek may possess unique strengths or optimizations that allow it to better adapt to the task’s complexities when competing with other models.